OpenStudy (anonymous):

Solve x + 3y = 9 3x - 3y = -13

OpenStudy (anteater):

You could use any of several methods, but in this case I think simple elimination would be easy. 3x - 3y = -13 x + 3y = 9 _______________ 4x + 0 = -4

OpenStudy (anteater):

And you can find x and y from there. :)